OpenStudy (anonymous):

lim as h ->0 of 1/h ln ((2+h)/2) is?

OpenStudy (raden):

ur question just similar to find derivative of f(x) = ln(2) do u know derivative of constanta will be .... ?

OpenStudy (anonymous):

use L'Hopitals Rule, if the limit is indeterminate, \[\lim_{x \rightarrow 0}\frac{ f(x) }{ g(x) } = \lim_{x \rightarrow 0}\frac{ f'(x) }{ g'(x) }\]so, \[\lim_{h \rightarrow 0}\frac{ \ln(\frac{ 2+h }{ 2 }) }{ h }=\lim_{h \rightarrow 0}\frac{ (\frac{ 2 }{ 2+h })(\frac{ 1 }{ 2 }) }{ 1 }\]\[=\frac{ 1 }{ 2 }\]